### v1

@@ 4.2 Varying the oversampling parameter $p$
The results are presented in Figure 2 and Table 2. In Figures 2a and 2b, AdaCUR was used with varying oversampling parameter $p$ on the parametric cookie problem (15). We set the tolerance $\epsilon$ to $10^{-12}$ and the error sample size $s$ to $5$ in this experiment. In Figure 2a, the algorithm meets the tolerance level up to a modest constant factor while approximately matching the $(\epsilon/\sqrt{n})$-rank in Figure 2b. The benefits of oversampling is demonstrated in Table 2 where the number of instances the algorithm recomputes the indices (which forms the dominant cost), $h_{2}$, decreases as the oversampling parameter $p$ increases. When $p=10$, the algorithm does not recompute the indices, and using only low-cost minor modifications, the algorithm rectifies the high relative error. That is, oversampling assists AdaCUR in improving its complexity by reducing the frequency of heavier computations.
In Figures 2c and 2d, we use FastAdaCUR. We vary the oversampling parameter $p$ on the parametric cookie problem (15) with the tolerance $\epsilon$ equal to $10^{-12}$ and the buffer size $b$ set to $5$. In Figure 2c, the algorithm fails to meet the tolerance $\epsilon=10^{-12}$ for certain parameter values, but still meets the tolerance within 2 orders of magnitude. By oversampling we are increasing the singular values of the core matrix $U_{i}=A(t_{i})(I,J)\in\mathbb{R}^{(r_{i}+b+p)\times(r_{i}+b)}$; see Section 3.3. This, in turn, should advocate a rank increase in FastAdaCUR, which the algorithm uses to reduce the error. However, as the algorithm is based on heuristics to prioritize speed, the impact that the parameters have on the algorithm is rather complicated and inconclusive. [^14] Running FastAdaCUR on the parametric cookie problem with the same parameter values as in Figure 2c several times yield different results with varying performance between different oversampling values, making the impact that oversampling has for the accuracy inconclusive. Nevertheless, for the accuracy and stability of the CUR decomposition, we recommend at least a little bit of oversampling; see [50].
(a) AdaCUR with $\epsilon=10^{-12}$, $s=5$
@@
Minor mod. only $h_{1}$ $15$ $7$ $8$
Recomput. Indices $h_{2}$ $5$ $1$ 00
@@ 4.3 Varying the error sample size $s$ and the buffer size $b$
The results are shown in Figure 3 and Table 3. In Figures 3a and 3b, we execute AdaCUR with varying error sample size $s$ on the Schrödinger equation (16). The oversampling parameter was set to $p=10$ and the tolerance $\epsilon=10^{-12}$. The algorithm satisfies the tolerance $\epsilon$ up to a modest constant factor as demonstrated in Figure 3a, while approximately aligning with the $(\epsilon/\sqrt{n})$-rank, as depicted in Figure 3b. As described in the role of the error sample size $s$ in Section 3.2, it makes the minor modifications in AdaCUR more effective. This is demonstrated in Table 3 where the number of times the algorithm recomputed the indices, $h_{2}$, decreases as the error sample size $s$ increases. The experiment demonstrates that with an increase in $s$, the algorithm becomes better at lowering the relative error using the low-cost minor modifications only. Therefore, the error sample size $s$ should be set sensibly to allow the algorithm to resolve the inaccuracy using minor modifications only so that the algorithm does not recompute indices until it becomes necessary. Note that higher $s$ increases the complexity of AdaCUR, so $s$ should not be too large.
In Figures 3c and 3d, we execute FastAdaCUR. We vary the buffer size $b$ on the Schrödinger equation 16 with the tolerance set to $\epsilon=10^{-12}$ and the oversampling parameter set to $p=10$. In Figure 3c, the algorithm performs well by meeting the tolerance except for the large spike happening in the first few parameter values. The cause of this spike is explained in Figure 3d where the problem experiences a substantial rank increase in the initial parameter values. Since the algorithm is only able to make a maximum of $b$ rank increase at each iteration, the algorithm struggles to keep up with the large rank increase in the initial parameter values. This observation is evident in Figure 3c, where the initial spike in the graph diminishes as $b$ increases. Therefore, if we expect a large rank increase in the parameter-dependent matrix, the buffer size $b$ should be set to a higher value for those parameter values.
@@
Err sample size $s$ or Buffer size $b$ $5$ $10$ $20$
Minor mod. only $h_{1}$ $16$ $13$ $11$
Recomput. Indices $h_{2}$ $20$ $5$ $2$
@@ 4.4 Adversarial example for FastAdaCUR
where $A_{1}=\mathtt{randn}(100,20)$, $A_{2}=\mathtt{randn}(200,10)$ and $c(t)=10^{-5+t/10}$ using the $\mathtt{randn}$ command in MATLAB. The parameter-dependent matrix $A(t)$ starts with a large component in the $A_{1}$ block, since $c(t)tA_{2}$ is the zero matrix at $t=0$. As $t$ increase, the dominant block becomes $c(t)tA_{2}$.
The results are presented in Figure 4. We execute FastAdaCUR in Figure 4b with tolerance $\epsilon=10^{-4}$ and various values of buffer size $b$ and oversampling parameter $p$. FastAdaCUR fails as the error continues to grow as we iterate through the parameter values. This can be explained by the nature of the algorithm. Since it only considers a submatrix of the original matrix for each parameter value, the $c(t)tA_{2}$ block remains hidden as $t$ increases. Consequently, the algorithm is never able to capture the $c(t)tA_{2}$ block, making the approximation poor. [^15] This type of counterexample is present in essentially all algorithms that do not view the entire matrix. For example, the algorithm in [17] may suffer from the same adversarial problem as it does not view the entire matrix at each parameter value. See [43, § 17.5] for a related discussion. On the other hand, as illustrated in Figure 4a, AdaCUR remains successful for the adversarial problem. AdaCUR is able to satisfy the tolerance of $10^{-4}$ and when the $c(t)tA_{2}$ block starts to become dominant, minor modifications or recomputation of indices is employed to capture the $c(t)tA_{2}$ block, making the algorithm successful.
@@
4.5 Speed test
When the target rank $k$ becomes sufficiently large, we anticipate that our algorithms, which run with $\mathcal{O}(T_{A(t)})$ and $\mathcal{O}((m+n)k^{2})$ complexity, will outperform many existing methods, which run with $\mathcal{O}(kT_{A(t)})$ complexity. We demonstrate this through experiments. Specifically, we compare AdaCUR and FastAdaCUR, against the randomized SVD and the generalized Nyström method from [37], as well as an algorithm in the dynamical low-rank approximation literature proposed by Lubich and Oseledets [41]. We use the following artificial problem. Let $A$ be a low-rank matrix with $\operatorname{rank}(A)=500$ given by $A=U\Sigma V^{T}\in\mathbb{R}^{50000\times 5000}$ where $U\in\mathbb{R}^{50000\times 500}$ and $V\in\mathbb{R}^{5000\times 500}$ are Haar-distributed orthonormal matrices and $\Sigma\in\mathbb{R}^{500\times 500}$ is a diagonal matrix with entries that decay geometrically from $1$ to $10^{-8}$. The parameter-dependent matrix $A(t)$ is then defined by
@@ 4.5 Speed test
where $i\in\{1,2,...,101\}$, $\delta=10^{-12}$ and $X_{j}$’s are sparse random matrices generated using $\mathtt{sprandn(50000,5000,10^{-5})}$ command in MATLAB. The parameter-dependent matrix $A(t)$ starts as a low-rank matrix at $i=1$, with sparse noise introduced at discrete time intervals. AdaCUR and FastAdaCUR’s input parameters were $\epsilon=10^{-6}$ for tolerance, $s,b=10$ for the error sample size and the buffer size, and the oversampling parameter was set to $p=10$. The target rank for the randomized SVD, generalized Nyström method and the algorithm by Lubich and Osedelets was set to the rank of $A$ which is $500$ with the second sketch size of the generalized Nyström set to $750$. See their respective papers for details [37, 41].
The results are illustrated in Figure 5. For AdaCUR and FastAdaCUR, we notice a slight rise at the beginning, stemming from the initial computation of indices. However, the low-rank approximation of the subsequent parameter values is computed very quickly, making the slope of the cumulative runtime graph relatively flat as we iterate through the parameter values. The graph is notably flat for FastAdaCUR, as it runs at most linearly in $m$ and $n$ at each iteration. On the other hand, the three existing methods display a steeper slope due to the higher computational cost associated with each parameter value. This higher cost stems from the increased number of matrix-vector products with $A(t)$ or its transpose. AdaCUR is approximately 44447777 times faster than the three existing algorithms, while FastAdaCUR is approximately 999916161616 times faster than the three existing algorithms.
@@ 5 Conclusion
In this work, we devised two efficient rank-adaptive algorithms for computing the low-rank approximation of parameter-dependent matrices: AdaCUR and FastAdaCUR. The key idea behind these algorithms is to try reuse the row and the column indices from the previous iterate as much as possible. AdaCUR comes with many favourable properties such as rank-adaptivity, error-control and a typical complexity of $\mathcal{O}(T_{A(t)})$, while FastAdaCUR, which is also rank-adaptive, is faster with a complexity that is at most linear in $m$ and $n$, but lacks error control, making it susceptible to adversarial problems. Nonetheless, FastAdaCUR should work on easier problems such as parameter-dependent matrices that are coherent at each parameter value.

### v2

@@ 4.2 Varying the oversampling parameter $p$
The results are presented in Figure 5 and Table 2. In Figures 5a and 5b, AdaCUR was used with varying oversampling parameter $p$ on the parametric cookie problem (16). We set the tolerance $\epsilon$ to $10^{-12}$ and the error sample size $s$ to $1$ in this experiment. In Figure 5a, the algorithm meets the tolerance level up to a modest constant factor while approximately matching the $(\epsilon/\sqrt{n})$-rank in Figure 5b. The benefits of oversampling is demonstrated in Table 2 where the number of instances the algorithm recomputes the indices (which forms the dominant cost), $h_{2}$, decreases as the oversampling parameter $p$ increases. This demonstrates that oversampling assists AdaCUR in improving its complexity by reducing the frequency of heavier computations.
In Figures 5c and 5d, we use FastAdaCUR. We vary the oversampling parameter $p$ on the parametric cookie problem (16) with the tolerance $\epsilon$ equal to $10^{-12}$ and the buffer size $b$ set to $5$. In Figure 5c, the algorithm fails to meet the tolerance $\epsilon=10^{-12}$ for certain parameter values, but still meets the tolerance within 2 orders of magnitude. By oversampling we are increasing the singular values of the core matrix $U_{i}=A(t_{i})(I,J)\in\mathbb{R}^{(r_{i}+b+p)\times(r_{i}+b)}$; see Section 3.3. This, in turn, should advocate a rank increase in FastAdaCUR, which the algorithm uses to reduce the error. However, as the algorithm is based on heuristics to prioritize speed, the impact that the parameters have on the algorithm is rather complicated and inconclusive. [^13] Running FastAdaCUR on the parametric cookie problem with the same parameter values as in Figure 5c several times yield different results with varying performance between different oversampling values, making the impact that oversampling has for the accuracy inconclusive. Nevertheless, for the accuracy and stability of the CUR decomposition, we recommend at least a little bit of oversampling; see [50].
(a) AdaCUR with $\epsilon=10^{-12}$, $s=1$
@@
Minor mod. only $h_{1}$ $24$ $18$ $5$
Recomput. Indices $h_{2}$ $6$ $5$ $3$
@@ 4.3 Varying the error sample size $s$ and the buffer size $b$
The results are shown in Figure 6 and Table 3. In Figures 6a and 6b, we execute AdaCUR with varying error sample size $s$ on the Schrödinger equation (17). The oversampling parameter was set to $p=10$ and the tolerance $\epsilon=10^{-12}$. The algorithm satisfies the tolerance $\epsilon$ up to a modest constant factor as demonstrated in Figure 6a, while approximately aligning with the $(\epsilon/\sqrt{n})$-rank, as depicted in Figure 6b. As described in the role of the error sample size $s$ in Section 3.2, it makes the minor modifications in AdaCUR more effective. This is demonstrated in Table 3 where the number of times the algorithm recomputed the indices, $h_{2}$, decreases as the error sample size $s$ increases. The experiment demonstrates that with an increase in $s$, the algorithm becomes better at lowering the relative error using the low-cost minor modifications only. Therefore, the error sample size $s$ should be set sensibly to allow the algorithm to resolve the inaccuracy using minor modifications only so that the algorithm does not recompute indices until it becomes necessary. Note that higher $s$ increases the complexity of AdaCUR, so $s$ should not be too large.
In Figures 6c and 6d, we execute FastAdaCUR. We vary the buffer size $b$ on the Schrödinger equation 17 with the tolerance set to $\epsilon=10^{-12}$ and the oversampling parameter set to $p=10$. In Figure 6c, the algorithm performs well by meeting the tolerance except for the large spike happening in the first few parameter values. The cause of this spike is explained in Figure 6d where the problem experiences a substantial rank increase in the initial parameter values. Since the algorithm is only able to make a maximum rank increase of $b$ at each iteration, the algorithm struggles to keep up with the large rank increase in the initial parameter values. This observation is evident in Figure 6c, where the initial spike in the graph diminishes as $b$ increases. Therefore, if we expect a large rank increase in the parameter-dependent matrix, the buffer size $b$ should be set to a higher value for those parameter values.
@@
Error sample size $s$ $5$ $10$ $20$
Minor mod. only $h_{1}$ $12$ $8$ $6$
Recomput. Indices $h_{2}$ $2$ $1$ 00
@@ 4.4 Adversarial example for FastAdaCUR
where $A_{1}=\mathtt{randn}(100,20)$, $A_{2}=\mathtt{randn}(200,10)$ and $c(t)=10^{-5+10t}$ using the $\mathtt{randn}$ command in MATLAB. The parameter-dependent matrix $A(t)$ starts with a large component in the $A_{1}$ block, since $c(t)tA_{2}$ is the zero matrix at $t=0$. As $t$ increases, the dominant block becomes $c(t)tA_{2}$.
The results are presented in Figure 7. We execute FastAdaCUR in Figure 7b with tolerance $\epsilon=10^{-4}$ and various values of buffer size $b$ and oversampling parameter $p$. FastAdaCUR fails as the error continues to grow as we iterate through the parameter values. This can be explained by the nature of the algorithm. Since it only considers a submatrix of the original matrix for each parameter value, the $c(t)tA_{2}$ block remains hidden as $t$ increases. Consequently, the algorithm is never able to capture the $c(t)tA_{2}$ block, making the approximation poor. [^14] This type of counterexample is present in essentially all algorithms that do not view the entire matrix. For example, the algorithm in [17] may suffer from the same adversarial problem as it does not view the entire matrix at each parameter value. See [43, § 17.5] for a related discussion. On the other hand, as illustrated in Figure 7a, AdaCUR remains successful for the adversarial problem. AdaCUR is able to satisfy the tolerance of $10^{-4}$, and when the $c(t)tA_{2}$ block starts to become dominant, minor modifications or recomputation of indices is employed to capture the $c(t)tA_{2}$ block, making the algorithm successful.
@@
4.5 Comparison against other methods
In this section, we test the data-driven aspect of AdaCUR and FastAdaCUR and compare their speed against other methods. Specifically, we benchmark AdaCUR and FastAdaCUR against the randomized SVD and the generalized Nyström method from [37], as well as an algorithm in the dynamical low-rank approximation literature proposed by Lubich and Oseledets [41].
4.5.1 Rank-adaptivity test
To test the rank-adaptive nature of AdaCUR and FastAdaCUR, we use the following synthetic problem. Let $A\in\mathbb{R}^{1000\times 300}$ be a low-rank matrix with $\operatorname{rank}(A)=50$, Haar-distributed left and right singular vectors, and singular values that decay geometrically from $1$ to $10^{-12}$. We define the parameter-dependent matrix $A(t)$ as
(19) $A(i)=A+10^{-4}\sum_{j=2}^{i}G_{1}^{(j)}G_{2}^{(j)}$
where $i\in\{1,2,...,10\}$, and $G_{1}^{(j)}\in\mathbb{R}^{1000\times 4}$, $G_{2}^{(j)}\in\mathbb{R}^{4\times 300}$ are Gaussian matrices with i.i.d. entries $\mathcal{N}(0,1/1000)$ and $\mathcal{N}(0,1/300)$ respectively. This experiment starts with a rank-$50$ matrix and incrementally adds rank-$4$ perturbations of size $10^{-4}$.
For AdaCUR and FastAdaCUR, the input parameters were set as follows: $\epsilon=10^{-9}$ for tolerance, $s,b=5$ for the error sample size and the buffer size, and $p=10$ for the oversampling parameter. For the randomized SVD, generalized Nyström method and the Lubich-Oseledets algorithm, the target rank was set to $50$, corresponding to the rank of $A$. The second sketch size of the generalized Nyström was set to $75$. See their respective papers [37, 41] for further details.
Figure 8: Rank-adaptivity test of AdaCUR and FastAdaCUR against the three existing algorithms: randomized SVD, generalized Nyström method [37] and the algorithm by Lubich and Osedelets [41].
The results are shown in Figure 8. We observe that fixed-rank methods–randomized SVD, generalized Nyström, and the Lubich-Oseledets algorithm–perform poorly as parameter value changes, failing to adapt to evolving data. In contrast, AdaCUR and FastAdaCUR effectively adjust the rank as needed to try maintain the error below the specified tolerance. This is guaranteed for AdaCUR with high probability due to its error control mechanism. This experiment highlights the data-driven nature of AdaCUR and FastAdaCUR, which dynamically adapt to the data, unlike the fixed-rank approaches. While AdaCUR and FastAdaCUR are effective rank-adaptive algorithms, they are not the only methods for addressing rank-adaptivity. Notably, rank-adaptivity has been explored in the dynamical low-rank approximation literature [8, 31, 32, 33]. This experiment highlights the simplicity and effectiveness of AdaCUR and FastAdaCUR in adapting to rank changes.
4.5.2 Speed test
When the target rank $r$ becomes sufficiently large, we anticipate that our algorithms, which run with $\mathcal{O}(T_{A(t)})$ and $\mathcal{O}((m+n)r^{2})$ complexity, will outperform many existing methods, which run with $\mathcal{O}(rT_{A(t)})$ complexity. We demonstrate this through experiments using the following artificial problem. Let $A$ be a low-rank matrix with $\operatorname{rank}(A)=500$ given by $A=U\Sigma V^{T}\in\mathbb{R}^{50000\times 5000}$ where $U\in\mathbb{R}^{50000\times 500}$ and $V\in\mathbb{R}^{5000\times 500}$ are Haar-distributed orthonormal matrices and $\Sigma\in\mathbb{R}^{500\times 500}$ is a diagonal matrix with entries that decay geometrically from $1$ to $10^{-8}$. The parameter-dependent matrix $A(t)$ is then defined by
@@ 4.5.2 Speed test
where $i\in\{1,2,...,101\}$, $\delta=10^{-12}$ and $X_{j}$’s are sparse random matrices generated using $\mathtt{sprandn(50000,5000,10^{-5})}$ command in MATLAB. The parameter-dependent matrix $A(t)$ starts as a low-rank matrix at $i=1$, with sparse noise introduced at discrete time intervals. AdaCUR and FastAdaCUR’s input parameters were $\epsilon=10^{-6}$ for tolerance, $s,b=10$ for the error sample size and the buffer size, and the oversampling parameter was set to $p=10$. The target rank for the randomized SVD, generalized Nyström method and the algorithm by Lubich and Osedelets was set to the rank of $A$, which is $500$. The second sketch size of the generalized Nyström was set to $750$. See their respective papers for details [37, 41].
The results are illustrated in Figure 9. For AdaCUR and FastAdaCUR, we notice a slight rise at the beginning, stemming from the initial computation of indices. However, the low-rank approximation of the subsequent parameter values is computed very quickly, making the slope of the cumulative runtime graph relatively flat as we iterate through the parameter values. The graph is notably flat for FastAdaCUR, as it runs at most linearly in $m$ and $n$ at each iteration. On the other hand, the three existing methods display a steeper slope due to the higher computational cost associated with each parameter value. This higher cost stems from the increased number of matrix-vector products with $A(t)$ or its transpose. AdaCUR is approximately 33335555 times faster than the three existing algorithms, while FastAdaCUR is approximately 777713131313 times faster than the three existing algorithms.
@@ 5 Conclusion
In this work, we devised two efficient rank-adaptive algorithms for computing the low-rank approximation of parameter-dependent matrices: AdaCUR and FastAdaCUR. The key idea behind these algorithms is to try to reuse the row and the column indices from the previous iterate as much as possible. AdaCUR comes with many favourable properties such as rank-adaptivity, error-control and a typical complexity of $\mathcal{O}(T_{A(t)})$, while FastAdaCUR, which is also rank-adaptive, is faster with a complexity that is at most linear in $m$ and $n$, but lacks error control, making it susceptible to adversarial problems. Nonetheless, FastAdaCUR should work on easier problems such as parameter-dependent matrices that are coherent at each parameter value.